### v2

# A.S.E: A Repository-Level Benchmark for Evaluating Security in AI-Generated Code

Abstract
The increasing adoption of large language models (LLMs) in software engineering necessitates rigorous security evaluation of their generated code. However, existing benchmarks often lack relevance to real-world AI programming scenarios, making them inadequate for assessing the practical security risks associated with AI-generated code in production environments. To address this gap, we introduce A.S.E (AI Code Generation Security Evaluation), a repository-level evaluation benchmark designed to closely mirror real-world AI programming tasks, offering a comprehensive and reliable framework for assessing the security of AI-generated code. Our evaluation of leading LLMs on A.S.E reveals several key findings. In particular, current LLMs still struggle with secure coding. The complexity in repository-level scenarios presents challenges for LLMs that typically perform well on snippet-level tasks. Morever, a larger reasoning budget does not necessarily lead to better code generation. These observations offer valuable insights into the current state of AI code generation, assisting developers in selecting the most appropriate models for practical tasks, while laying the foundation for refining LLMs to generate secure and efficient code in real-world applications.
22footnotetext: Co-first authors.11footnotetext: Corresponding authors: Hui Li (lih64@pkusz.edu.cn) and Dong Zhang (zalezhang@tencent.com).
1 Introduction
The rapid advancement of large language models (LLMs) has greatly enhanced the AI programming ecosystem, with tools like Cursor and Claude Code enabling developers to choose models that best fit their tasks. These AI assistants significantly improve programming efficiency, leading to a surge of AI-generated code in production environments. However, research @8SecurityEval ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; fu2024constraineddecodingsecurecode has shown that such code can harbor security vulnerabilities, posing serious risks such as data breaches or system failures. Relying on developers to ensure the security of AI-generated code can be highly challenging given the complexity of modern software systems. Therefore, there is a pressing need to identify and utilize AI models that are capable of generating secure code.
Despite the numerous benchmarks developed by both academia and industry to evaluate AI-generated code, most of them  DBLP:journals/corr/abs-2107-03374@17HumanEval ; DBLP:journals/corr/abs-2108-07732@19MBPP ; DBLP:conf/iclr/0003XM24@28RepoBench ; DBLP:journals/corr/abs-2406-11612@31Long-code-arena ; DBLP:conf/nips/DingWADTJRNBRX23@29CrossCodeEval ; DBLP:conf/acl/LiangJH025@32REPOCOD ; DBLP:conf/acl/LiZGM0PHWL25@33FEA-Bench ; jimenez2024swebench primarily focus on code quality, such as syntax correctness and functional accuracy, while overlooking critical security considerations. Although some benchmarks @8SecurityEval ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; fu2024constraineddecodingsecurecode attempt to address code security (as shown in Table 1), they are often inadequate to assess the actual security risks of AI-generated code in real-world production scenarios due to several key reasons: (a) Limited relevance to real-world data. These datasets are typically sourced from human-curated synthetic code snippets, which have limited relevance to the functional and security scenarios of real-world projects. (b) Code generation tasks detached from real-world AI programming. Their code generation tasks are typically limited to isolated code snippets, focusing solely on functional descriptions without considering the context within files or projects, which does not align with mainstream AI programming paradigm. (c) Unreliable code evaluation methods. Security assessments of generated code often rely on manual or LLM-based judgment, which are unreliable and difficult to automate or reproduce consistently. This gap poses a significant challenge for both developers and organizations seeking to integrate AI-generated code securely into their systems.
To bridge this gap, we introduce A.S.E (AI Code Generation Security Evaluation), a repository-level evaluation benchmark designed to closely mirror real-world AI programming scenarios, offering a comprehensive and reliable framework for assessing the security of AI-generated code. Specifically, A.S.E has the following key design features: (a) Real-world data source: The dataset is derived from high-quality GitHub open-source repositories with documented CVEs. A.S.E leverages vulnerability-related code extracted from CVE patches, ensuring that the data reflects both realistic and security-sensitive scenarios. (b) Simulation of real-world code generation tasks: A.S.E mimics AI programming assistants like Cursor by extracting code contexts (including both intra-file and cross-file contexts) from the repository, and providing them to LLMs for code generation. (c) High accuracy and reproducibility in code evaluation: For each test case, corresponding to a specific CVE and repository, A.S.E designs targeted static vulnerability detection rules that can scan for the original CVE, ensuring accurate security assessment of the regenerated project.
Building on these design principles, the A.S.E benchmark includes 120 repository-level instances, consisting of 40 seed datasets collected from GitHub and 80 mutated variants generated through semantic and structural mutation techniques, such as identifier renaming and control-flow reshaping. These variants are introduced to mitigate data leakage risks, ensuring that the evaluation reflects the LLM’s capabilities rather than its memorization. To closely simulate real-world code generation processes, A.S.E uses the BM25 algorithm to automatically extract relevant code context from the repository when designing code generation tasks. These code generation tasks cover four common vulnerability types in real-world web projects. We focus on the web domain due to the primary application of AI programming in web development in production environments. These vulnerability-related codes to be generated are closely tied to the projects’ business logic, requiring models to understand both project code and security considerations. A.S.E also incorporates five commonly used web development languages to assess how programming languages influence model performance. Furthermore, A.S.E integrates customized SAST rules for security evaluation on each data instance in the dataset. While A.S.E aims to evaluate the security of AI-generated code, it also considers code quality and stability. Given that security is meaningful only when code correctness is ensured, and in light of potential hallucinations in large model outputs, A.S.E uses multi-dimensional metrics that assess code security, quality, and generation stability for a comprehensive evaluation of the model’s capabilities.
Based on A.S.E, we evaluated 26 mainstream commercial and open-source models under the same experimental setup, leading to several key findings. First, existing LLMs still face significant challenges in secure coding. All models fall short in terms of security performance compared to their code quality performance (such as syntax correctness). Even the best-performing model, Claude-3.7-Sonnet, achieved only a total score of 52.79 in our evaluation. Second, A.S.E introduces significant complexity in repository-level scenarios, which presents a challenge for LLMs that typically perform well on snippet-level tasks. For example, although GPT-3 excels on SafeGenBench DBLP:journals/corr/abs-2506-05692@13SafeGenBench , its performance on the A.S.E. benchmark drops, falling behind many other models. Third, slow-thinking configurations, which allocate more deliberate computation or multi-step reflection, tend to underperform compared to fast-thinking configurations that rely on concise, direct decoding. This suggests that a larger reasoning budget does not necessarily lead to better code generation. These observations offer valuable insights into the current state of AI code generation, helping developers select the most appropriate models for their specific tasks. Furthermore, they provide a foundation for refining LLMs, enhancing their ability to generate secure and efficient code in real-world applications.
The main contributions of this paper are summarized as follows:
  • New repository-level benchmark from real code. We release A.S.E, a repository-level evaluation benchmark derived from real-world GitHub repositories with documented CVEs. Unlike existing benchmarks, A.S.E is designed to closely mirror real-world AI programming tasks by leveraging vulnerability-related code from CVE patches, ensuring the data reflects both realistic and security-sensitive scenarios.
  • Automated and reproducible evaluation framework. We develop a reproducible vulnerability-targeted evaluation framework that integrates custom vulnerability detection rules tailored to each data instance. Compared to previous work, this framework enables more automated and accurate code evaluation. It comprehensively considers the capabilities of AI-generated code, including security, quality, and generation stability.
  • Extensive experiments and findings. We evaluate 26 mainstream commercial and open-source LLMs on A.S.E, revealing several key findings. These insights shed light on the current state of AI code generation, guiding developers in selecting the most suitable models for their tasks. Additionally, they lay the groundwork for refining LLMs to improve their ability to generate secure and efficient code in real-world applications.
2 Related Work
In recent years, numerous benchmarks have been developed for evaluating AI-generated code, which can be broadly categorized into two types. Functionality-oriented benchmarks DBLP:journals/corr/abs-2107-03374@17HumanEval ; DBLP:journals/corr/abs-2108-07732@19MBPP ; DBLP:conf/iclr/0003XM24@28RepoBench ; DBLP:journals/corr/abs-2406-11612@31Long-code-arena ; DBLP:conf/nips/DingWADTJRNBRX23@29CrossCodeEval ; DBLP:conf/acl/LiangJH025@32REPOCOD ; DBLP:conf/acl/LiZGM0PHWL25@33FEA-Bench ; jimenez2024swebench –the majority of existing efforts–primarily assess whether the generated code is syntactically correct and functionally accurate, with limited emphasis on security or vulnerability detection. For example, HumanEval DBLP:journals/corr/abs-2107-03374@17HumanEval evaluates models through algorithmic-contest programming tasks to measure functional correctness in language understanding, algorithms, and basic mathematics. Security-oriented benchmarks @8SecurityEval ; DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench ; fu2024constraineddecodingsecurecode , on the other hand, go a step further by evaluating whether the generated code is secure and reliable. Table 1 summarizes representative benchmarks in this category and compares them with A.S.E. In what follows, we highlight the distinctions between A.S.E and existing benchmarks from two perspectives: relevance to real-world scenarios and code assessment methods.
Table 1: Comparison of Security-Oriented Code Generation Evaluation Datasets
Dataset
CWE
Tags
Granularity
(Repo/Snippet)
Provenance
Domain
Open
Source
Security
Eval.
[…]
SafeGenBench Snippet Human-Curated Synthetic Task-Based Coding Problems SAST + LLM
BaxBench Snippet Human-Curated Synthetic Backend Programming Tasks Test Cases
CWEval Snippet Human-Curated Synthetic Task-Based (CWE-Tagged) Test Cases
CODEGUARD+ Snippet Human-Curated Synthetic Task-Based Coding Problems Test Cases
CodeLMSec Snippet Human-Curated Synthetic Task-Based Coding Problems SAST
SecurityEval Snippet Human-Curated Synthetic Task-Based Coding Problems SAST + Manual
2.1 Relevance to real-world scenarios
Early functionality-oriented benchmarks for code generation typically emphasized small, self-contained tasks at the function or snippet level (e.g., HumanEval (DBLP:journals/corr/abs-2107-03374@17HumanEval, ), MBPP (DBLP:journals/corr/abs-2108-07732@19MBPP, )), using unit tests or reference outputs to measure basic functional correctness. Their datasets are often constructed from human-curated synthetic sources, such as algorithm competition problems, which provide well-defined but limited scenarios. Due to their simplicity, standard code benchmarks have quickly become saturated. Consequently, more challenging benchmarks with stronger relevance to real-world scenarios have emerged DBLP:conf/iclr/0003XM24@28RepoBench ; DBLP:journals/corr/abs-2406-11612@31Long-code-arena ; DBLP:conf/nips/DingWADTJRNBRX23@29CrossCodeEval ; DBLP:conf/acl/LiangJH025@32REPOCOD ; DBLP:conf/acl/LiZGM0PHWL25@33FEA-Bench ; jimenez2024swebench . For example, SWE-Bench jimenez2024swebench evaluates the code generation capabilities of LLMs and agents by framing tasks around issue resolution in real-world projects.
In contrast, security-oriented benchmarks remain at a relatively early stage. Their scope is still limited, with most focusing on snippet-level tasks that lack strong connections to real-world programming scenarios (@8SecurityEval, ; DBLP:journals/corr/abs-2502-11844@12BaxBench, ; DBLP:conf/llm4code/PengCHYR25@11CWEval, ; DBLP:conf/satml/HajipourHHSF24@9CodeLMSec, ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench, ; fu2024constraineddecodingsecurecode, ). This restricts their ability to capture the complexity of security issues in practical development environments. To address this gap, we introduce A.S.E, which, similar to SWE-Bench, is derived from real-world projects. Unlike existing security-oriented benchmarks, A.S.E builds on repositories with documented CVEs and incorporates repository-level context into the code generation process, thereby ensuring that evaluations are both realistic and security-sensitive.
2.2 Code Assessment Methods.
Prior security-oriented benchmarks have adopted diverse methods for assessing the security of AI-generated code, yet each comes with notable limitations. (a) Manual evaluation offers flexibility but is labor-intensive and difficult to scale. For example, SecurityEval @8SecurityEval relies on expert validation, which makes it difficult to automate and scale to new model evaluations. (b) LLM-as-judge approaches scale more easily and capture semantic nuances, but their judgments are highly sensitive to prompt design, model version, and decoding randomness, making results unreliable and hard to reproduce (DBLP:journals/corr/abs-2312-04724@25CyberSecEval, ). (c) Generic static analysis tools (SAST) provide deterministic rule-based detection, yet often suffer from false positives and false negatives when applied across different languages or without contextual calibration DBLP:conf/satml/HajipourHHSF24@9CodeLMSec ; DBLP:journals/corr/abs-2506-05692@13SafeGenBench . (d) Test-case-based assessments are straightforward but inherently limited DBLP:journals/corr/abs-2502-11844@12BaxBench ; DBLP:conf/llm4code/PengCHYR25@11CWEval ; fu2024constraineddecodingsecurecode . For instance, BaxBench DBLP:journals/corr/abs-2502-11844@12BaxBench designs correctness and security tests that are agnostic to frameworks and programming languages, but this generality makes it difficult to fully validate generated code, leading to under-reporting of vulnerabilities. To overcome these challenges, A.S.E emphasizes reproducibility and precision through customized vulnerability detection. For each data instance—corresponding to a specific project and CVE—A.S.E designs tailored static analysis rules that explicitly model the relevant sources, sinks, and taint propagation paths. These rules are calibrated to accurately detect the original vulnerability within the unmodified repository, ensuring that the evaluation framework can reliably assess whether regenerated code reintroduces or mitigates the same flaw. By automating this process at the project level, A.S.E achieves more reliable, scalable, and reproducible security assessments compared to prior approaches.
3 The A.S.E Benchmark
This section introduces the A.S.E. framework, which includes three core components: benchmark construction (subsection 3.2), code generation task setup (subsection 3.3), and code evaluation (subsection 3.4), as shown in Figure 1. We will first highlight the key features of the A.S.E. design, followed by a detailed discussion of each of these three core components.
Figure 1: Overall workflow of A.S.E. (a) A.S.E benchmark construction: from high-quality GitHub seeds, we build the A.S.E dataset via dual mutations (structure/semantic), CVE patches, and a customized SAST tool, followed by expert curation. (b) Model code generation: given an incomplete repository, a vulnerability description and context guide LLMs to complete the repository. (c) Security evaluation: comprehensive assessment with security, quality and stability.
3.1 Design Philosophy
We aim to create a repository-level evaluation benchmark that mirrors real-world AI programming scenarios, offering a reliable framework for assessing the security of AI-generated code. To ensure accurate results, A.S.E. focuses on realistic data sources, task settings, and code assessment methods. Specifically, the core features of the A.S.E benchmark include the following three components:
(i) Data Source: Real-world and Repository-Level Data Sources. To reflect the performance of large models in real-world software environments, A.S.E constructs tasks from active open-source repositories with documented CVEs and verifiable patches. It utilizes vulnerability-related code extracted from CVE patches, ensuring the data captures realistic and security-sensitive scenarios. To mitigate the risk of data leakage, A.S.E employs semantic and structural mutation techniques, such as identifier renaming and control-flow reshaping, on the collected real-world repositories. These variants help prevent data leakage and ensure that the evaluation reflects the LLM’s capabilities, not its memorization.
(ii) Task Settings: Real-world AI Programming Task Settings. To simulate realistic usage, A.S.E replicates AI programming assistants like Cursor by extracting code contexts—both intra-file and cross-file—directly from the repository. These contexts are then provided to LLMs for code generation, closely mimicking real-world AI programming scenarios. Moreover, the generated code is output in the form of diff files, allowing patches to be applied directly to the repository, further reflecting real software development practices.
(iii) Code Assessment: High Accuracy and Reproducibility Assessment. Instead of relying on manual or LLM-based judgment, which can be unreliable and difficult to reproduce consistently, A.S.E designs targeted static vulnerability detection rules for each test case, corresponding to a specific CVE and repository. These rules are tailored with dedicated source–sink definitions and taint propagation paths to successfully detect the original CVE, thereby ensuring an accurate security assessment of the regenerated project.
Following these guidelines, we introduce the A.S.E benchmark to evaluate in repository-level code generation regarding security. Next, we introduce the three key steps: the construction of the benchmark, the design of the code generation task, and the evaluation of the results.
Figure 2: Overview of A.S.E benchmark construction. Algorithm-guided screening and preselection (left): aggregate CVE-linked sources and automatically filter repositories by web-related CWEs, vulnerability types (XSS, SQL injection, path traversal, command injection), and languages (Java, Python, Go, PHP, JavaScript). Expert-guided curation and refinement (right): conduct manual review, reproducibility and exploitability checks, and dual-toolchain SAST (e.g., CodeQL + Joern) with CVE-specific rules; then expand $40$ seed repositories via structural/semantic mutation to $80$ variants.
3.2 A.S.E Benchmark Construction
Guided by our design philosophy, we construct the A.S.E benchmark as shown in Figure 1 (a) and 2. To ensure realistic scenarios and relevant security expertise, we organize a dedicated team of ten contributors with strong backgrounds in cybersecurity and web development from top-tier universities, including five Ph.D. candidates and five master-level students. All contributors have extensive hands-on experience in vulnerability discovery, analysis, and remediation, with emphasis on common web vulnerabilities such as XSS, SQL injection, and path traversal. Each contributor adheres to strict secure coding standards and is familiar with static code analysis techniques. The construction of the benchmark proceeds in four stages: determining data sources, filtering candidate repository, expert-guided refinement and quality filtering, and dataset expansion. The following subsections will provide a detailed explanation of each step.
Step 1: Determining Data Sources. We collect real-world software projects from both enterprise-internal vulnerability repositories and public platforms like GitHub. This stage ensures representativeness and diversity, so that the dataset covers a broad range of software types and vulnerability scenarios.
Step 2: Filtering candidate repository. Then we apply strict criteria to filter candidate repositories, ensuring both project quality and vulnerability completeness. Only repositories with active maintenance, sufficient test coverage, and consistent development practices are retained. In addition, each repository must contain at least one publicly disclosed CVE accompanied by a complete patch diff at the commit level. For every vulnerability case, detailed metadata—such as a textual description, CVE identifier, severity rating, and reproducible steps—must be available. These conditions guarantee that the selected projects provide realistic software contexts and verifiable vulnerability information, thereby ensuring high data quality and reproducibility for subsequent stages.
Step 3: Expert-Guided Refinement and Quality Control. To ensure that the collected data is security-relevant and represents real-world vulnerabilities that are both detectable and reproducible, we refine the dataset further through expert-guided analysis and systematic filtering. After the initial screening, security experts examine each candidate repository to precisely annotate vulnerable code regions. The system then extracts surrounding context—including function signatures, API definitions, and call chains—to reconstruct the code environment. For every vulnerability, we design CodeQL or Joern queries and verify the corresponding taint propagation paths. Projects with ambiguous or inconsistent results, or with code that current tools cannot analyze with sufficient precision, are discarded. For the retained cases, detection rules are customized and refined to improve accuracy and ensure alignment with confirmed vulnerabilities. Once validated, the labeled vulnerable code is removed to create a fill-in-the-code setting. By combining the functional description of the vulnerability with its extracted context, we generate structured prompts that allow models to reason over repository-level structures and logic rather than isolated snippets. After expert review, 40 repositories with verified CVE records are selected as the seed dataset, each with a baseline commit serving as a stable starting point for task construction and evaluation. This process guarantees authenticity, reliability, and reproducibility across all benchmark tasks.
Step 4: Dataset Expansion. After constructing base tasks from real-world vulnerability repositories, we expand the dataset to increase both volume and coverage under a semantics-preserving constraint. Two categories of transformations are applied. The first is semantic transformations, which diversify surface expressions by systematically renaming variables and functions or substituting equivalent APIs. The second is structural transformations, which alter control flow, refactor call graphs, or reorganize file layouts to introduce structural variation. These operations preserve functional behavior and code semantics while modifying implementation details, thereby reducing overlap with public repository code that may appear in model training corpora and mitigating potential data contamination. The expanded variants enable a more comprehensive assessment of model robustness and generalization.
Figure 3: Statistics of A.S.E benchmark.
3.2.1 General Statistics
After the construction, we obtain the A.S.E. benchmark, which comprises $120$ repository-level vulnerability instances, including $40$ seed repositories and $80$ mutated variants.
Specifically, the benchmark targets four vulnerability categories that are common in real web projects. Each category maps one-to-one to a corresponding CWE entry: SQL Injection ($29.2\%$, CWE-89), Path Traversal ($26.7\%$, CWE-22), Cross-Site Scripting ($25.0\%$, CWE-79), and Command Injection ($19.2\%$, CWE-78). This mapping defines the dataset at the CWE level and limits it to these four categories.
From the perspective of programming languages, A.S.E encompasses five mainstream environments that reflect realistic multi-language software development. The distribution is concentrated in PHP ($50.0\%$), followed by Python ($19.2\%$), Go ($14.2\%$), JavaScript ($14.2\%$), and Java ($2.5\%$). This distribution highlights the prevalence of vulnerable PHP code in real-world applications while also offering opportunities to explore cross-language generalization in evaluation.
3.3 Model Code Generation
After the construction, we proceed to set up tasks for large language models (LLMs) based on the collected data. The primary goal is to assess the model’s ability to generate code within a real-world repository context. To achieve this, we design tasks that involve filling in vulnerable code regions with contextually appropriate completions.
The task setup begins by removing the labeled vulnerable code from each repository, creating a "fill-in-the-code" environment. We then combine the functional description of the vulnerability with the extracted context to generate a structured code-completion prompt for each task. This setup ensures that the input-output semantics are clear, allowing the model to reason over the repository’s structure and logic, rather than merely generating local code snippets. As a result, the evaluation focuses on the model’s ability to understand context and generate code accordingly.
Specifically, for each benchmark instance, A.S.E. first retrieves the corresponding GitHub repository and checks out the baseline commit containing the vulnerability. Expert annotations are used to automatically mask the vulnerable region in the target file, which is replaced by the special token <masked>. The LLM then receives an input combining two key components: (i) the masked file with a functional description generated by Claude-Sonnet-4 anthropic2025systemcard@claude-4 and refined by experts, and (ii) repository-level context that includes related source files selected by BM25 DBLP:journals/ftir/RobertsonZ09@bm25 ranking and the README of the project.
To ensure that the generated code can be directly applied to the repository for evaluation, models are instructed to produce output in patch format, which is automatically applied using commands such as git apply. We ensure reproducibility by packaging each scenario with Docker and running each instance three times under identical conditions.
3.4 Code Assessment
After the models generate code, we evaluate the outputs along multiple dimensions including quality, security, and stability. Given that security is meaningful only when code is functionally correct, we first perform a pre-check for quality. This check ensures that the generated patch can be successfully applied to the repository and passes basic static checks, such as syntax verification. Next, we assess security by measuring whether the applied patch reduces the number of detected vulnerabilities, using expert-crafted static analysis rules tailored to each benchmark instance. Finally, to account for the inherent variability of large language models, we introduce a stability metric, which evaluates the consistency of model outputs across repeated runs for the same task. This provides insight into how reliably a model generates correct and secure code.
Next, we detail the evaluation metrics used in the pipeline. The following definitions formalize these steps.
Quality. This measures whether a generated patch successfully integrates and passes essential static checks. A test is considered successful only if the patch merges cleanly and satisfies both static analysis and syntax checks. The quality score is then defined as:
$\text{Quality}=\frac{1}{N}\sum_{t=1}^{N}q_{t},$ (1)
where $N$ denotes the total number of tests, and $q_{t}=1$ if test $t$ merges and passes all checks, and $q_{t}=0$ otherwise.
Security. This measures the effectiveness of a generated patch in reducing vulnerabilities. Specifically, A.S.E. applies expert-crafted static analysis rules tailored to each CVE to count vulnerabilities before and after patch integration. The security score for each model is then computed as:
$\text{Security}=\frac{1}{N}\sum_{t=1}^{N}s_{t},$ (2)
where $s_{t}=1$ if $v_{\text{after}}(t)<v_{\text{before}}(t)$ and $s_{t}=0$ otherwise, and where $v_{\text{before}}(t)$ and $v_{\text{after}}(t)$ denote the numbers of detected vulnerabilities before and after patch integration for test $t$.
Stability. This measures the consistency of model’s generated code across repeated runs for the same benchmark instance. Higher stability indicates that the model produces more predictable and reliable outputs. To quantify this, we first compute the standard deviation of a model’s results over three independent runs for each benchmark instance $i\in\mathcal{B}$, denoted as $\sigma_{i}$. To convert lower variation into a higher score, we normalize the values using min-max scaling:
$\tilde{\sigma}_{i}=\begin{cases}1-\dfrac{\sigma_{i}-\sigma_{\min}}{\sigma_{\max}-\sigma_{\min}},&\text{if }\sigma_{\max}>\sigma_{\min},\\ 1,&\text{otherwise},\end{cases}$ (3)
where $\sigma_{\min}=\min_{i}\sigma_{i}$ and $\sigma_{\max}=\max_{i}\sigma_{i}$. The stability score is computed as follows:
$\text{Stability}=\frac{1}{\lvert\mathcal{B}\rvert}\sum_{i\in\mathcal{B}}\tilde{\sigma}_{i}$ (4)
The overall score combines the three dimensions with fixed weights as follows:
[…]
This pipeline enables reproducible project level testing and provides a systematic assessment of functional correctness and security robustness of patches generated by large language models in realistic development settings.
4 Experiments
In this section, we evaluate a representative set of large language models on the A.S.E. benchmark to examine their ability in generating secure code. We first describe the evaluated models, then present overall results, and finally conduct detailed analyses to derive key findings on the strengths and limitations of current models.
Evaluated Models. We choose a total of 26 state-of-the-art large language models (LLMs), consisting of 18 proprietary models and 8 open-source models. A key selection criterion is the availability of both “fast thinking” and “slow thinking” modes, which allows for a comprehensive comparison of reasoning paradigms.
For the proprietary models, our evaluation covers flagship systems across multiple families. This includes the Claude series (Claude-3.7-Sonnet (@claude35S, ), Claude-Sonnet-4 (anthropic2025systemcard@claude-4, ), Claude-Opus-4 (anthropic2025systemcard@claude-4, )) and their “thinking” counterparts, the GPT family (GPT-4o (DBLP:journals/corr/abs-2410-21276, ), GPT-4.1 (openai_gpt4.1_2025, ), Codex-mini (openai_codexmini_2025, ), and additional variants), the Grok series (Grok-3 (xai_grok3_2025, ), Grok-4 (xai_grok4_2025, ), Grok-3-mini (xai_grok3_2025, )), Gemini-2.5-Pro (DBLP:journals/corr/abs-2507-06261, ), Qwen-Coder-Plus (hui2024qwen2.5-coder, ), and Hunyuan-T1 (tencent_hunyuanT1_2025, ).
For the open-source models, we select 8 widely adopted representatives spanning diverse architectures. The set includes the Qwen3 series (DBLP:journals/corr/abs-2505-09388, ) (Qwen3-235B-A22B-Instruct, Qwen3-Coder, Qwen3-235B-A22B), DeepSeek-V3 (DBLP:journals/corr/abs-2412-19437, ), DeepSeek-R1 (DBLP:journals/corr/abs-2501-12948, ), Kimi-K2 (team2025kimi, ), and GLM-4.5zeng2025glm .
4.1 Overall Results
As shown in 2, we evaluated 26 state-of-the-art LLMs on the A.S.E. benchmark. The table reports each model’s scores across three dimensions: Code Security, Code Quality, and Generation Stability. The “License” column indicates whether a model is proprietary or open-source, while the “Thinking” column denotes its reasoning mode (fast-thinking or slow-thinking). Models are ranked in descending order based on their overall scores.
Overall, the results reveal a substantial gap between code quality and security: while most models produce syntactically correct and useful code, none surpass the 50-point threshold on Code Security. This indicates that secure coding remains a critical weakness for current LLMs. The A.S.E. benchmark effectively exposes LLMs’ weaknesses in secure code generation. As a repository-level evaluation, it requires cross-file dependency resolution and long-context reasoning, going beyond isolated snippet-level generation. Consequently, models that perform well on snippet-oriented benchmarks, such as GPT-o3 on SafeGenBench DBLP:journals/corr/abs-2506-05692@13SafeGenBench , experience a significant drop in performance.
Among the evaluated models, Claude-3.7-Sonnet achieves the highest overall score ($63.01$) and a strong Code Quality score ($91.58$), yet its Code Security score remains below 47. Similarly, Claude-Sonnet-4 obtains the best Code Quality performance ($92.37$) but only $34.78$ in Code Security. In contrast, GPT-o3 exhibits extremely high Generation Stability (98.91) but fail almost completely in security and quality. Similar patterns appear in GPT-4.1 and Qwen3-235B-A22B. This indicates that high stability does not guarantee secure code.
Moreover, Figure 4 presents the distribution of code generation outcomes for each model, categorized into four types: qualified and secure, qualified but insecure, patch integration failed, and SAST check failed. These results highlight two main patterns: (1) Flagship models tend to prioritize code correctness over security. For example, Claude-3.7-Sonnet generates 91.7% qualified code, yet 43.8% of it remains insecure. (2) Weaker models struggle with basic code generation in complex repository-level scenarios, producing a lower proportion of qualified code and failing most SAST checks.
Table 2: The leaderboard of various advanced Code LLMs on the A.S.E. benchmark. [Uncaptioned image]is the fast-thinking mode and [Uncaptioned image]indicates slow-thinking mode.
Rank Model License Thinking Overall Security Quality Stability
1 Claude-3.7-Sonnet-20250219 Proprietary [Uncaptioned image] 63.01 46.72 91.58 75.00
2 Claude-3.7-Sonnet-Thinking-20250219 Proprietary [Uncaptioned image] 61.04 44.65 89.85 72.92
3 Qwen3-235B-A22B-Instruct-2507 Open Source [Uncaptioned image] 60.15 48.03 82.08 67.08
4 Qwen3-Coder Open Source [Uncaptioned image] 59.31 42.69 85.16 81.54
5 DeepSeek-V3-20250324 Open Source [Uncaptioned image] 58.59 40.89 85.87 82.94
6 Claude-Sonnet-4-20250514 Proprietary [Uncaptioned image] 57.14 34.78 92.37 85.65
7 Kimi-K2-20250711-Preview Open Source [Uncaptioned image] 55.29 37.82 79.90 86.25
8 GPT-4o-20241120 Proprietary [Uncaptioned image] 55.10 45.65 72.46 59.67
9 Qwen-Coder-Plus-20241106 Proprietary [Uncaptioned image] 53.55 37.98 73.78 86.27
10 Claude-Opus-4-20250514 Proprietary [Uncaptioned image] 52.71 31.95 85.82 77.91
11 Grok-3 Proprietary [Uncaptioned image] 52.18 38.64 73.54 69.41
12 DeepSeek-R1-20250528 Open Source [Uncaptioned image] 51.76 38.01 74.39 66.38
13 Gemini-2.5-Pro-Exp-20250325 Proprietary [Uncaptioned image] 51.02 29.98 84.04 78.21
14 Claude-Sonnet-4-Thinking-20250514 Proprietary [Uncaptioned image] 50.92 34.10 76.81 74.22
15 Claude-Opus-4-Thinking-20250514 Proprietary [Uncaptioned image] 50.17 30.70 79.84 77.98
16 GLM-4.5 Open Source [Uncaptioned image] 49.80 35.92 70.24 71.74
17 Grok-4 Proprietary [Uncaptioned image] 42.40 29.53 59.78 67.42
18 o4-mini-20250416 Proprietary [Uncaptioned image] 41.35 27.87 60.74 64.07
19 Grok-3-mini Proprietary [Uncaptioned image] 30.49 22.37 38.15 56.26
20 Codex-mini-latest Proprietary [Uncaptioned image] 29.71 22.96 34.68 55.29
21 Hunyuan-T1-20250321 Proprietary [Uncaptioned image] 21.92 15.57 20.21 65.18
22 Qwen3-235B-A22B-Thinking Open Source [Uncaptioned image] 18.11 9.42 15.60 77.81
23 GPT-4.1-20250414 Proprietary [Uncaptioned image] 17.26 5.26 16.46 91.66
24 Qwen3-235B-A22B Open Source [Uncaptioned image] 13.37 3.34 7.27 91.86
25 o3-mini-20250131 Proprietary [Uncaptioned image] 13.23 3.67 3.91 98.57
26 o3-20250416 Proprietary [Uncaptioned image] 10.22 0.36 0.36 98.91
Figure 4: Attributional distribution across Code LLMs.
4.2 Detailed Analysis and Findings
In the following analysis, we examine model performance from multiple complementary perspectives to uncover systematic patterns in security and code quality. We start by comparing proprietary and open-source models, followed by an investigation of architectural and scaling effects. Next, we assess the impact of reasoning paradigms (fast vs. slow thinking), and then break down performance at the task level. Finally, we illustrates representative error modes with case studies.
I. Model category: Open-source models perform comparably to closed-source Code LLMs. Our results reveal that the performance gap between leading open-source and closed-source Code LLMs is narrowing. While top-tier closed-source models like Claude-3.7-Sonnet and Claude-Sonnet-4 achieve marginally higher Code Quality scores ($91.58$ and $92.37$, respectively), prominent open-source models such as Kimi-K2 and Qwen-Coder-Plus demonstrate superior generation stability. In terms of overall performance, the competition is remarkably tight. For instance, the open-source model Qwen3-235B-A22B-Instruct ranks second, closely following the leading Claude-3.7-Sonnet series. This outcome indicates that the performance gap between open-source and closed-source Code LLMs is narrow.
II. Reasoning paradigms: Slow-thinking Paradigms can lead to security regressions. We found that reasoning paradigms designed for more deliberate and careful generation ("slow-thinking") tend to underperform in Code Security compared to their "fast-thinking" counterparts. As illustrated in 5, this trend is consistent across multiple models. For example, Claude-3.7-Sonnet-Thinking scores $44.65$ in Code Security, a slight degradation from its fast-thinking counterpart. The drop is even more pronounced for Claude-Sonnet-4-Thinking, which lags its non-thinking version across all metrics. This pattern suggests that while slow-thinking aims to improve reasoning, it may inadvertently increase security risks, possibly by generating more complex code or lacking targeted security reinforcement.
Figure 5: Fast v.s. Slow Thinking.
III. Model Architecture: MoE models generally outperform dense models. Although the architectures of some closed-source models remain undisclosed, nearly all leading open-source Code LLMs adopt a Mixture-of-Experts (MoE) architecture, including Qwen3-235B-A22B, Qwen3-Coder-480B-A35B-Instruct, DeepSeek-V3-671B-A37B, and Kimi-K2-Preview-1T-32B. This trend indicates that MoE-based Code LLMs generally achieve stronger security performance than dense models.
IV. Task-level challenges: Path Traversal Presents the Greatest Challenge. To assess model performance across different vulnerability types, we analyzed Claude-3.7-Sonnet, one of the better-performing models. As shown in 6, Path Traversal is consistently the most challenging task. Among the four evaluated vulnerability types, all Code LLMs perform relatively weakly on Path Traversal, with even the most advanced model scoring below $50.0$. This difficulty likely stems from the subtlety and context-dependence of path manipulation techniques, which are harder to detect than more explicit attacks. The results suggest that current Code LLMs lack robust reasoning about file system operations and access control. Enhancing model understanding of file path construction and improving generalization across diverse traversal scenarios are essential for stronger defense against this vulnerability.
Figure 6: Detailed performance of various Code LLMs across four tasks of A.S.E benchmark.
Figure 7: Detailed Attribution classification of Claude-3.7-Sonnet: Original (top) and Mutation Test (bottom).
V. Scaling Law on Code Security. Beyond cross-model comparisons, an open question is whether security performance scales with model size. To address this, we evaluate the Qwen2.5-Coder-Instruct and Qwen3 series. As shown in Table 3, larger parameter scales generally yield better performance, particularly in Overall and Quality. Within the Qwen3-Instruct series, a clear scaling trend on Security emerges, improving from 33.57 $\rightarrow$ 45.46 $\rightarrow$ 48.03 as model size increases. In contrast, the Qwen2.5-Coder-Instruct series shows growth that eventually plateaus with scale. Beyond this, results further indicate that Qwen3 consistently outperforms Qwen2.5-Coder across all metrics, reflecting stronger architectural and training advantages; meanwhile, Instruct variants achieve substantially higher scores than their Thinking counterparts.
VI. Benchmark Consistency Across Original and Mutated Datasets. We further examine whether models exhibit performance differences between the original benchmark and its mutated variant. Across the evaluated models, the results show minimal variation before and after mutation, which suggests that the benchmark is robust, free from substantial data leakage, and effective in manual construction. For illustration, Fig. 7 presents the attribution classification of a representative model, Claude-3.7-Sonnet, across different vulnerability types (original test at the top, mutation test at the bottom). As we can observe, for Path Traversal and SQL Injection, the model frequently produces code that is well-formed yet insecure, whereas for XSS and Command Injection, it more often generates outputs that are both secure and well-qualified. Failures such as patch integration errors or SAST check violations are rare, indicating that the model generally succeeds in producing correct and usable code at the repository level. However, a considerable fraction of outputs still contain latent security vulnerabilities. This consistency across datasets confirms the benchmark’s robustness, while underscoring that secure code generation remains a persistent challenge.
Table 3: Qwen model performance by scale. Bold numbers indicate the best score per series.
Model Overall Security Quality Stability
Qwen2.5-Coder Series
0.5B-Instruct 36.67 25.56 37.79 100.00
1.5B-Instruct 31.57 26.86 32.53 56.90
3B-Instruct 34.12 29.52 38.28 49.22
7B-Instruct 45.60 40.78 52.95 52.47
14B-Instruct 42.76 32.24 56.44 64.87
32B-Instruct 44.43 30.99 65.08 63.16
Qwen3 Series
4B-Thinking-2507 39.93 33.57 44.43 64.57
4B-Instruct-2507 39.05 32.08 49.17 50.50
30B-A3B-Thinking-2507 41.89 31.85 56.21 59.20
30B-A3B-Instruct-2507 56.59 45.46 72.89 74.47
235B-A22B-Thinking-2507 35.18 24.51 46.89 64.09
235B-A22B-Instruct-2507 60.15 48.03 82.08 67.08
5 Conclusion
This paper addressed the critical need for realistic evaluation of LLM-generated code security, as existing benchmarks often use isolated snippets and lack reproducible, context-aware methods. We introduced A.S.E (AI Code Generation Security Evaluation), a benchmark built from real-world repositories with documented CVEs. By preserving full project context and using a reproducible, containerized framework with expert-defined rules, A.S.E reliably assesses code security, build quality, and stability, directly linking them to the model’s input. Our evaluation of leading LLMs yielded three key findings: (1) While Claude-3.7-Sonnet leads in overall performance, (2) the security gap to open-source models is narrow, with Qwen3-235B-A22B-Instruct achieving the top security score. (3) Crucially, concise "fast-thinking" decoding strategies consistently outperform complex "slow-thinking" reasoning for security repairs. These results have significant implications, suggesting prompting strategy is as critical as model choice and that open-source models are highly competitive in security. A.S.E provides a robust foundation for future research into developing more secure and reliable LLMs for software engineering.
6 Data Availability
To facilitate further research and practical adoption, we have made both the source code of the A.S.E evaluation framework and the full benchmark dataset publicly available, which can be accessed at https://github.com/Tencent/AICGSecEval. In addition, we provide comprehensive evaluation results for all tested models, which are continuously updated and maintained on the project website https://aicgseceval.tencent.com/home. By releasing both the framework and dataset, we aim to support reproducibility, encourage community contributions, and promote the development of more secure and reliable AI-assisted programming.